OpenStudy (anonymous):

Solve by the substitution method.... 9m+n=45 m-5n=51

OpenStudy (anonymous):

Rewrite 45m+5n=225 m-5n = 51 SUbstitude 2 from 1 44m = 174 m = 174/44 n = (174/44 -51)/5

OpenStudy (anonymous):

so what is the order pair then please

OpenStudy (anonymous):

This does no make scence to me Ivanbaj 45m+5n=225 how did you get that?

OpenStudy (anonymous):

Problem! You need to add 1. To 2 !!! 46m = 276 m =276/46

OpenStudy (anonymous):

m = 6 n= 9

OpenStudy (anonymous):

(6,9)

OpenStudy (amistre64):

m-5n=51 <- m = 51+5n 9(51+5n)+n=45 ; solve for n 459 + 45n +n = 45 46n = -414 n = -414/46 = -9 ............

OpenStudy (amistre64):

m = 51+5n m = 51+5(-9) m = 51 - 45 m = 6

OpenStudy (cherrilyn):

9m+n=45 m-5n=51 ..get m alone to get m=5n+51. then plug that into the first equation 9(5n+51)=45 45n+459=45 get n alone.. 45n=-414 divide both sides by 45 n=-9.2 plug n into 9m+n=45 to get m so 9m+(-9.2)=45 9m=54.2 divide both sides by 9 so you get m=6.02 and n=-9.2

OpenStudy (cherrilyn):

oops its 9(5n+51)+n=45 45n+459+n=45 46n=-414 n=-9 plug n = -9 into 9m+(-9)=45 9m=54 m=6